### v1

@@ 3 Discussion
Despite advancements in GDL and the growing interest in aggregation-based models for functional brain connectome analysis [23, 25, 4, 24, 46, 3, 26, 27, 28], our findings suggest that sophisticated GDL models do not necessarily outperform simpler models, such as classical ML models and MLPs. This challenges the prevailing assumption that increased model complexity leads to better predictive performance.
@@
Many existing works implicitly assume that GDL’s complexity guarantees superior results. In practice, these studies often omit direct comparisons with simpler models [23, 25, 4, 28], thereby leaving the true effectiveness of these complex models inadequately assessed. Even when simpler baselines are included, as in BrainGNN [3] and NeuroGraph [26], benchmarking is incomplete, and their experimental conclusions are inconsistent. For instance, BrainGNN shows SVM and Random Forest outperforming graph models in some cases, while MLPs underperform; conversely, NeuroGraph finds MLPs significantly outperforming most graph models, with weaker Random Forest results. Our study address the limitations from these works by rigorously optimizing and benchmarking both simple and complex models, providing a more comprehensive evaluation of various models for brain connectome analysis.
@@
The trend towards developing increasingly complex models to capture intricate patterns in brain data, should be approached with caution. Our results suggest that simple models, when properly tuned, can serve as strong baselines and, in some cases, even outperform more complex models. This highlights the need for a more balanced perspective where simple models are not only included as baselines but also carefully evaluated to ensure that any proposed advancements genuinely offer superior performance.
Ineffectiveness of Message Aggregation
Message aggregation techniques are widely adopted across domains to capture complex relational information [36, 38, 37], but in functional brain networks, our results show that these operations degrade rather than improve prediction accuracy. GDL-based brain connectome analysis models typically use the connection profile as the node features, which has become a standard practice [23, 25, 4, 24, 3, 26, 27, 28, 30]. Specifically, the connection profile, representing a row of the functional brain connectome matrix, inherently captures the global relationships between a given node and all other nodes in the network. Since these node features already contain the full spectrum of inter-node connections, the aggregation employed in GDL models [29], which pools information from neighboring nodes, introduces redundancy and even destroy originally obvious features. This is related to over-smoothing, a phenomenon where repeated aggregation causes node features to become indistinguishable from one another [47], ultimately diluting the discriminative power of the original features. Moreover, connection profiles are derived from BOLD time series correlations, reflecting strong linear relationships between ROIs, which also leads to redundancy. Consequently, complex models are more prone to over-smoothing during aggregation, making it difficult to extract meaningful high-order features. In contrast, simpler models like logistic regression better utilize these linear relationships and avoid the potential noise from aggregation.
Additionally, the nature of functional connectivity plays an important role. Connections in functional brain networks are typically based on statistical correlations rather than true causal or direct neuronal interactions [8, 7, 48]. Aggregation in GDL assumes valid information pathways between connected nodes [49, 50], which may not hold in functional brain networks. Consequently, the model tends to integrate statistically correlated but causally or structurally irrelevant noise into the updates of node features, which diminishes predictive performance.
Performance of the Dual-pathway Model
@@
The proposed hybrid model, which integrates both linear and graph-based components, achieves robust prediction performance across datasets. The LM pathway leverages the strong predictive power inherent in the original functional brain connectivity, preserving the direct input-output relationship and avoiding feature distortions. Meanwhile, the GAT pathway improves upon previous models by using BOLD time series embeddings as node features, instead of connection profiles. BOLD time series embeddings retain the raw temporal dynamics and variability of ROIs’ signals, which captures richer temporal information and mitigates over-smoothing issues linked to the strong linear dependencies and redundancy in connection profiles.
Interpretability of the Dual-pathway Model
@@
Important Connections (Edge-level): The dual-pathway model provides complementary insights into brain network organization, highlighting both localized and global patterns of connectivity for understanding the neural basis of brain function. The GAT pathway’s focus on within-system connections, particularly within DMN, SM and CE, resonates with well-established findings in neuroscience about cognitive ability. The DMN is tied to internally directed cognitive processes such as self-referential thinking, memory retrieval, and future planning [51, 52, 53]. Similarly, the SM network, crucial for motor control and sensory processing, has shown within-network coherence related to fine motor skills and their correlation with cognitive abilities [13, 54]. The CE network’s within-system emphasis reinforces its critical role in executive functions like working memory, decision-making, and cognitive control, which is correlated with fluid intelligence [55, 56]. This within-system focus reflects the specialized, localized processing within functional networks. Moreover, the observed bilateral symmetry suggests that the GAT pathway effectively captures the functional coherence between homologous regions across hemispheres, which is known to be important for cognitive stability and performance [14, 57]. On the other hand, the identification of cross-system connections between DMN-CE [58, 9], as well as between the Vis-SM systems [10, 11], underscores the importance of network integration in cognitive function.
The LM pathway, on the other hand, captures a more diffusive and global connectivity pattern, balancing positive and negative weights across brain networks. This balance indicates LM pathway’s ability to account for both facilitative and inhibitory interactions across ROIs, necessary for maintaining cognitive flexibility [12, 17]. The dispersed weighting also reflects the integration of long-range connections necessary for large-scale network communication and distributed processing across the brain functional systems [2, 59, 60]. Overall, the model’s ability to capture the dual aspects of brain connectivity offers a more comprehensive interpretation of the neural mechanisms underlying cognitive processes, bridging the gap between localized processing and global brain network integration.
Salient ROIs (Node-level): The concentration of attention-based ROIs within the CE and DMN aligns well with their well-documented roles in cognitive control and introspective processes. The involvement of the Orbital, Polar, and Inferior Frontal cortexes aligns with their contributions to executive functions and decision-making [61, 62, 63]. The identification of key regions within the Medial Temporal cortex by the GAT pathway also aligns with its known role in memory formation and retrieval [64, 15], critical for reasoning and knowledge application. The engagement of both frontal and temporal lobes highlights the distributed nature of cognitive processes [65] underlying fluid intelligence.
@@
The LM pathway adds further interpretability through its positive and negative weights. Positive weights, indicating regions where increased activity correlates with higher fluid intelligence, are prominent within the VS and SM systems. The VS, particularly involving the Insular and Frontal Opercular cortexes, is crucial for detecting and prioritizing relevant stimuli [66], while the SM supports motor planning and cognitive execution [13]. In contrast, negative weights, indicating that increased activity is associated with lower fluid intelligence, are found mainly in DMN, especially within the Orbital and Polar Frontal cortexes. This aligns with the “Default Mode Interference Hypothesis” [67], positing that excessive DMN activity, associated with mind-wandering and self-referential thought, can interfere with task-focused cognition.
@@
Graph Properties of the Highlighted Brain Connectivity (Subgraph-level): In the GAT-induced subgraph, the skewed degree distribution reflects the scale-free nature typical of biological networks. Such networks are known for their resilience to random failures, which supports the brain’s robustness in maintaining cognitive functions [2, 68]. The higher clustering coefficient and modularity suggest that the attention mechanism captures tightly-knit communities, consistent with the brain’s modular organization, where distinct networks support specialized functions [8, 43]. The notably higher small-worldness metric further highlights a balance between local specialization and global integration [69, 70], a hallmark of efficient brain organization.
Conversely, the LM-induced subgraph, with shorter average path lengths and higher global efficiency, suggests a network optimized for rapid information transfer and efficient brain communication [71, 72]. The broader degree distribution indicates a more even network [65] without dominant hubs, possibly reflecting distributed processing across brain regions. Higher assortativity coefficient suggests a preference for connections between similar nodes [8], consistent with local homogeneity in certain brain areas [73, 74].
@@
In summary, the dual-pathway model offers a comprehensive view of brain connectivity, with each pathway contributing distinct interpretative information. The GAT pathway excels at capturing localized, modular structures, emphasizing key functional hubs and the brain’s small-world properties. In contrast, the LM pathway highlights global integration and network efficiency, focusing on distributed processing and long-range connections. Together, the proposed hybrid model provides a balanced view, bridging specialized modular functions with broader network communication, thus deepening our understanding of the neural basis of brain functions.
Limitations and Future Extensions
In this study, we extensively evaluate aggregation-based GDL models for functional brain connectome analysis across multiple datasets and tasks. While these models do not outperform simpler models in prediction, they offer potential in interpretability. Besides these findings, several limitations should be acknowledged to provide a comprehensive view of this research:
@@
Modality Focus: We exclusively utilize rs-fMRI data to construct functional connectivity networks. However, the brain’s structural connectivity, often derived from modalities like Diffusion Tensor Imaging (DTI), provides complementary information about the tangible physical pathways between ROIs. Integrating structural connectivity might offer a more holistic understanding of the role of GDL models for brain connectome analysis.
Node Features Choice: This study emphasizes the use of connection profiles and BOLD time-series data as node features, given their prevalence in current research. Nevertheless, other node feature representations, such as graph theoretical metrics or neurological contextual information specific to each ROI– such as regional brain function, structural properties, or known involvement in particular cognitive or clinical processes, could also be valuable in brain connectome analysis.
Graph Construction Methods: The experiments in this study rely on correlations to construct the functional brain networks. While correlation is a widely accepted measure of functional connectivity, alternative graph construction methods, such as those based on causal inference or adaptive graph construction during the learning process, are not explored. These methods could potentially capture different aspects of brain connectivity by accounting for confounding effects or by generating task-specific connections in an end-to-end manner, which might influence the effectiveness of aggregation-based models.
@@ 4 Methods
  • Autism Brain Imaging Data Exchange (ABIDE) [31]: a publicly available dataset [31] consists of anonymized resting-state functional magnetic resonance imaging (rs-fMRI) data collected from 17 international sites, providing a diverse and comprehensive collection for studying Autism Spectrum Disorder (ASD). The data is preprocessed using the C-PAC pipeline [75], which includes several stages: slice time correction using AFNI’s 3dTshift, motion correction with AFNI’s 3dvolreg, skull-stripping with AFNI’s 3dAutomask, global mean intensity normalization and band-pass filtering (0.01-0.1 Hz). After quality control, 1009 subjects are left, with 516 individuals (51.14%) diagnosed with ASD. The prediction task focuses on binary classification for ASD diagnosis. The brain regions are defined using the Craddock 200 atlas [76], with each subject represented by a $200\times 200$ functional connectivity matrix. The BOLD time-series for each ROI are truncated to a length of 100.
@@
  • Philadelphia Neuroimaging Cohort (PNC) [32]: a dataset originates from the Brain Behavior Laboratory at the University of Pennsylvania and the Children’s Hospital of Philadelphia. We preprocess this rs-fMRI dataset through a series of steps including slice timing correction, motion correction, registration, normalization, removal of linear trends, bandpass filtering (0.01-0.1 Hz), and spatial smoothing. After preprocessing and quality control, the dataset includes 504 subjects, with 289 (57.34%) being female. The prediction task involves binary classification for gender prediction. The regions of interest are parcellated using the 264-node atlas defined by Power et al. [77]. Each subject has time-series data collected through 120 time steps.
@@
  • The Human Connectome Project (HCP) [33]: a systematic effort to map macroscopic human brain circuits and their relationship to behavior in a large population of healthy adults. For our analysis, we utilize resting-state fMRI data from the HCP dataset, processed using the standard pipeline in the CONN toolbox [78], involving realignment, co-registration, normalization, and smoothing of the raw images. Confounding effects from motion artifacts, white matter, and cerebrospinal fluid (CSF) are regressed out of the signal. Subsequently, the BOLD time series are band-pass filtered (0.01-0.1 Hz) to retain relevant functional connectivity information. We focus on the 132 nodes defined on the Harvard-Oxford atlas [79]. After rigorous quality control, this dataset includes 982 samples. Each sample is represented by a $132\times 132$ functional connectivity matrix. The prediction task is fluid intelligence score prediction, treated as a regression problem, with the scores having a mean of 17.03 and a standard deviation of 4.70. To ensure consistency, the BOLD time-series are truncated to a length of 1024.
@@
  • The Adolescent Brain Cognitive Development Study (ABCD) [34]: a large-scale, longitudinal project that recruits children aged 9-10 years across 21 sites in the United States. This study includes repeated imaging scans, along with extensive psychological and cognitive assessments. For our analysis, we use rs-fMRI data from the baseline visit, which is processed using the standard, open-source ABCD-HCP BIDS fMRI pipeline [^1] https://github.com/DCAN-Labs/abcd-hcp-pipeline. The preprocessing pipeline includes multiple stages, starting with distortion correction, anatomical alignment, and surface reconstruction. Further, the functional images undergo motion correction, distortion removal and registration. ANTs-based bias field correction and global signal regression are used to enhance data quality. The time series are band-pass filtered between 0.008 and 0.09 Hz. After quality control, we get 7,327 samples, each represented by a $360\times 360$ connectivity matrix, constructed by calculating the Pearson correlation between the BOLD time-series of each pair of ROIs. The regions of interest (ROIs) are defined based on the HCP 360 atlas [80]. The prediction task is fluid intelligence score prediction with a mean of 18.21 and a standard deviation of 3.69, treated as a regression problem. Considering the variability in sequence lengths within the ABCD dataset, the BOLD time-series are truncated to a unified length of 512 for consistency.
We intentionally select a diverse set of datasets with varying preprocessing procedures, ROI parcellations, and prediction tasks to demonstrate the robustness and generalizability of our findings. This diversity reduces the likelihood of coincidental results and enhances the applicability of our conclusions to a broader range of brain connectome analysis studies.
Benchmarked baselines and the proposed dual-pathway model
@@
Classical Machine Learning Baselines: Connectome-based Predictive Modeling (CPM) [19], a widely recognized model in network neuroscience for predicting behavioral or clinical outcomes based on brain connectivity data. CPM is typically split into two variations: CPM-NEG and CPM-POS, where predictive features are selected based on negative and positive correlations with the target variable, respectively. Covariance Maximizing Eigenvector-based Predictive Modeling (CMEP) [22], a more recent model that first generates candidate features based on eigenvectors with strong linear covariance with the target variable. These features are then used in ElasticNet regression to assess their predictive power, streamlining the process by avoiding arbitrary feature selection thresholds. Additionally, we employed classical models including Logistic Regression for classification, Linear Regression for regression, ElasticNet for both classification and regression, Support Vector Machine (SVM) for classification, Support Vector Regression (SVR) for regression, Random Forest for both classification and regression, as well as Naive Bayes for classification and Kernel Ridge Regression for regression. An overview of these models can be found in [39].
@@
These models are chosen for their established use in brain connectome analysis, ensuring that our findings could be directly compared and validated against existing literature. For all models except CMEP, the input features are the vectorized upper triangular part of the functional brain network matrix, a setting commonly used in classical ML-based functional brain network analysis studies [17, 12, 18, 19, 20, 21]. CMEP uses the entire symmetric functional brain network matrix as input.
@@
Most of these classical ML models require a feature selection process. For CPM, feature selection follows the original methodology by applying a p-value threshold [19] to the correlation between each connection (feature) and the target variable. Features with p-values below the threshold are retained for model training and prediction. CMEP, however, does not require explicit feature selection as it inherently identifies predictive features during its eigenvector generation process. For the other classical ML models, feature selection involves selecting the top $m$ features with the smallest p-values. Specifically, for classification tasks (e.g., using the ABIDE and PNC datasets), p-values are determined through a t-test, while for regression tasks, they are calculated based on the correlation between features and the target variable.
@@
Non-graph Deep Learning Baselines: BrainNetCNN [81], a convolutional neural network (CNN) framework specifically designed for predicting clinical neurodevelopmental outcomes from brain networks. Unlike classical image-based CNNs that use spatially local convolutions, BrainNetCNN incorporates novel convolutional filters that operate on the topological locality of structural brain networks. These filters include edge-to-edge, edge-to-node, and node-to-graph convolutions, making the model particularly well-suited for leveraging the unique structure of brain connectivity data. Additionally, we also use two multi-layer perceptron (MLP)-based models: MLP-Flatten and MLP-Node. MLP-Flatten takes the upper triangular part of the brain connectome matrix as input, flattens it into a feature vector, and processes it through two fully connected layers. In contrast, MLP-Node transforms each node’s feature individually using shared weights across nodes, then concatenates these transformed node features into a graph-level embedding for the final prediction.
Aggregation-based Graph Deep Learning Baselines: Message aggregation in GDL refers to combining information from neighboring nodes to update the representation of a central node, which is crucial in learning the structural and relational patterns within graph data. Below, we describe the GDL models employed in our study, focusing on general-purpose GDL techniques and those specifically tailored for brain connectome analysis.
@@
The general-purpose GDL baselines include following models: Graph Convolutional Network (GCN) [40], a fundamental graph neural network (GNN) model that generalizes convolutional operations to graph structures. It aggregates the features of neighboring nodes by applying a convolution-like operation, which computes a weighted sum of the neighboring node features. Graph Attention Network (GAT) [35] introduces an attention mechanism, allowing the model to assign different importance to neighboring nodes. The attention coefficients are adaptively learned during training, enabling the network to focus on the most relevant neighbors. This adaptive aggregation strategy makes GAT particularly useful for capturing the heterogeneity in node connections, where certain connections might be more relevant to the task. Graph Isomorphism Network (GIN) [41] is designed to be as powerful as the Weisfeiler-Lehman graph isomorphism test. It aggregates neighboring node features by applying a multi-layer perceptron (MLP) to the sum of the features. GraphSAGE [42] updates node embeddings by averaging the features of neighboring nodes and passing the result through an MLP.

### v2

@@ 3 Discussion
Despite advancements in GDL and the growing interest in aggregation-based models for functional brain connectome analysis [braingb, fbnetgen, bnt, ibgnn, dag, braingnn, neurograph, gated, community_bnt], our findings suggest that sophisticated GDL models do not necessarily outperform simpler models, such as classical ML models and MLPs.
@@
Many existing works implicitly assume that GDL’s complexity guarantees superior results. In practice, these studies often omit direct comparisons with simpler models [braingb, fbnetgen, bnt, community_bnt], thereby leaving the true effectiveness of these complex models inadequately assessed. Even when simpler baselines are included, as in BrainGNN [braingnn] and NeuroGraph [neurograph], benchmarking is incomplete, and their experimental conclusions are inconsistent. For instance, BrainGNN shows SVM and Random Forest outperforming graph models in some cases, while MLPs underperform; conversely, NeuroGraph finds MLPs significantly outperforming most graph models, with weaker Random Forest results. Our study address the limitations from these works by rigorously optimizing and benchmarking both simple and complex models, providing a more comprehensive evaluation of various models for brain connectome analysis.
@@
Our results suggest that simple models, when properly tuned, can serve as strong baselines and, in some cases, even outperform more complex models. This highlights the need for a more balanced perspective where simple models are not only included as baselines but also carefully evaluated to ensure that any proposed advancements genuinely offer superior performance.
Although message aggregation techniques are widely adopted across domains to capture complex relational information [social_recommentation, he2020lightgcn, molecular], in functional brain networks, our results show that these operations degrade rather than improve prediction accuracy. GDL-based brain connectome analysis models typically use the connection profile as the node feature, which represents one row of the functional connectivity matrix and inherently encodes the correlations between a given region of interest (ROI) and all others [braingb, fbnetgen, bnt, ibgnn, braingnn, neurograph, gated, community_bnt, hypergraph]. While standard in the field, this design implies that each node feature already captures global network information and shows considerable redundancy and inter-node correlation even before any aggregation.
This intrinsic redundancy is quantified in Table. 1, where the effective rank of the input node-feature matrices is found to be substantially lower than the feature dimensionality (7–38% of feature dimensions across datasets), indicating that the input features reside in a low-dimensional subspace. In such a low-rank regime, message aggregation, which essentially acts as a low-pass filter [li2018deeper] over the graph, tends to further amplify the shared components of node features and suppress distinctive variations, thereby progressively diminishing the predictive capability of the learned node representations, a process consistent with the over-smoothing [smoothing] phenomenon observed in GDL models.
These observations suggest a mismatch between the prevailing formulation of functional brain connectome analysis (Pearson correlation–based connectivity with connection profile node features) and the inductive bias of message aggregation in current GDL models. Improving the alignment between brain representation and model architecture may help enhance the effectiveness of GDLs applied to functional brain connectome analysis. Future studies could explore alternative graph formulations and model variants that better account for the redundancy and low-rank characteristics of functional connectivity data, aiming to capture more meaningful and predictive neural patterns.
@@
Motivated by these observations, we develop a dual-pathway hybrid model that jointly leverages linear and graph-based representations to capture complementary aspects of brain connectivity. This hybrid model achieves robust predictive performance across datasets. The LM pathway leverages the strong predictive power inherent in the original functional brain connectivity, preserving the direct input-output relationship and avoiding feature distortions. Meanwhile, the GAT pathway improves upon previous models by using BOLD time series embeddings as node features, instead of connection profiles. BOLD time series embeddings retain the raw temporal dynamics and variability of ROIs’ signals, which captures richer temporal information and mitigates over-smoothing issues linked to the strong dependencies and redundancy in connection profiles.
To further understand how the dual-pathway model captures functional brain organization, we examined its interpretability from three complementary perspectives: edge-, node-, and subgraph-level patterns.
@@
At the edge level, the GAT pathway’s focus on within-system connections, particularly within DMN, SM and CE, resonates with well-established findings in neuroscience about cognitive ability. The DMN is tied to internally directed cognitive processes such as self-referential thinking, memory retrieval, and future planning [self_ref1, self_ref3, mem_ret3]. Similarly, the SM network, crucial for sensory processing, has shown within-network coherence related to cognitive abilities [sm1, sm2]. The CE network’s within-system emphasis reinforces its critical role in executive functions like working memory, decision-making, and cognitive control, which is correlated with fluid intelligence [ce2, ce4]. This within-system focus reflects the specialized, localized processing within functional networks. Moreover, the observed bilateral symmetry suggests that the GAT pathway effectively captures the functional coherence between homologous regions across hemispheres, which is known to be important for cognitive stability and performance [sym2, sym3]. While the identification of cross-system connections between DMN-CE [dmn_ce1, dmn_ce2], as well as between the Vis-SM systems [vim_sm2, vim_sm3], underscores the importance of network integration in cognitive function. The LM pathway, on the other hand, captures a more diffusive and global connectivity pattern, balancing positive and negative weights across brain networks. This balance indicates LM pathway’s ability to account for both facilitative and inhibitory interactions across ROIs, necessary for maintaining cognitive flexibility [pos_neg2, pos_neg3]. The dispersed weighting also reflects the integration of long-range connections necessary for large-scale network communication and distributed processing across the brain functional systems [long_range1, long_range2, long_range3].
@@
At the node level, the model highlights salient regions of interest (ROIs) that align well with established neural substrates of cognitive control and introspective processing. The concentration of attention-based ROIs within the CE and DMN aligns well with their well-documented roles in cognitive control and introspective processes. The involvement of the Orbital, Polar, and Inferior Frontal cortexes aligns with their contributions to executive functions and decision-making [orbital_frontal1, orbital_frontal3, orbital_frontal4]. The identification of key regions within the Medial Temporal cortex by the GAT pathway also aligns with its known role in memory formation and retrieval [medial_temp1, medial_temp2], critical for reasoning and knowledge application. The engagement of both frontal and temporal lobes highlights the distributed nature of cognitive processes [dist2] underlying fluid intelligence. The LM pathway adds further interpretability through its positive and negative weights. Positive weights, indicating regions where increased activity correlates with higher fluid intelligence, are prominent within the VS and SM systems. The VS, particularly involving the Insular and Frontal Opercular cortexes, is crucial for detecting and prioritizing relevant stimuli [insular_frontal], while the SM supports motor planning and cognitive execution [sm1]. In contrast, negative weights, indicating that increased activity is associated with lower fluid intelligence, are found mainly in DMN, especially within the Orbital and Polar Frontal cortexes. This aligns with the “Default Mode Interference Hypothesis” [hypothesis], positing that excessive DMN activity, associated with mind-wandering and self-referential thought, can interfere with task-focused cognition.
@@
At the subgraph level, the GAT-induced subgraph, the skewed degree distribution reflects the scale-free nature typical of biological networks. Such networks are known for their resilience to random failures, which supports the brain’s robustness in maintaining cognitive functions [long_range1, scale_free1]. The higher clustering coefficient and modularity suggest that the attention mechanism captures tightly-knit communities, consistent with the brain’s modular organization, where distinct networks support specialized functions [complex_measures, clustering_coef1]. The notably higher small-worldness metric further highlights a balance between local specialization and global integration [small_world1, small_world3], a hallmark of efficient brain organization. Conversely, the LM-induced subgraph, with shorter average path lengths and higher global efficiency, suggests a network optimized for rapid information transfer and efficient brain communication [communication2, communication3]. The broader degree distribution indicates a more even network [dist2] without dominant hubs, possibly reflecting distributed processing across brain regions. Higher assortativity coefficient suggests a preference for connections between similar nodes [complex_measures], consistent with local homogeneity in certain brain areas [homogeneity1, homogeneity2].
@@
In summary, the dual-pathway model offers a comprehensive view of brain connectivity, with each pathway contributing distinct interpretative information. The GAT pathway excels at capturing localized, modular structures, emphasizing key functional hubs and the brain’s small-world properties. In contrast, the LM pathway highlights global integration and network efficiency, focusing on distributed processing and long-range connections. Together, the proposed hybrid model provides a balanced view, bridging specialized modular functions with broader network communication, thus deepening our understanding of the neural basis of brain functions. It is important to note that the interpretability analysis on the ABCD dataset is presented as a representative case study illustrating the dual-pathway model’s ability to capture complementary local and global neural organization patterns, rather than to derive age-specific or cohort-dependent neuroscientific conclusions.
@@
Finally, several limitations of this study should be acknowledged, along with potential avenues for future extensions. First, We exclusively utilize rs-fMRI data to construct functional connectivity networks. However, the brain’s structural connectivity, often derived from modalities like Diffusion Tensor Imaging (DTI), provides complementary information about the tangible physical pathways between ROIs. Integrating structural connectivity might offer a more holistic understanding of the role of GDL models for brain connectome analysis.
Second, in this study, we focus on the mainstream setting in which functional brain networks are constructed using Pearson correlation, following the most widely adopted practice in fMRI-based connectome analysis. All major observations and conclusions in this work are derived under this standard setting. Nonetheless, there still exist many other approaches to mapping functional connectivity [liu2025benchmarking], such as distance-based, more sophisticated causality-based, or spectral methods, as well as end-to-end adaptive graph learning strategies that infer connectivity during model training [fbnetgen], which we have not yet explored. These alternative formulations might influence the effectiveness of GDL models in distinct ways and represent promising directions for future investigation.
Third, this study primarily evaluates GDLs using two node-feature families: connection profiles and BOLD time-series embeddings. Results and conclusions are therefore grounded in these choices. Other options—e.g., graph-theoretical metrics, additional BOLD-derived handcrafted features, and ROI-specific neurological context—remain unexplored and may differentially affect GDL behavior and interpretability, motivating future studies in this direction.
@@ 4 Methods
The Autism Brain Imaging Data Exchange (ABIDE) dataset [abide] is a publicly available dataset consists of anonymized resting-state functional magnetic resonance imaging (rs-fMRI) data collected from 17 international sites, providing a diverse and comprehensive collection for studying Autism Spectrum Disorder (ASD). The data is preprocessed using the C-PAC pipeline [cpac], which includes several stages: slice time correction using AFNI’s 3dTshift, motion correction with AFNI’s 3dvolreg, skull-stripping with AFNI’s 3dAutomask, global mean intensity normalization and band-pass filtering (0.01-0.1 Hz). After quality control, 1009 subjects are left, with 516 individuals (51.14%) diagnosed with ASD. The prediction task focuses on binary classification for ASD diagnosis. The brain regions are defined using the Craddock 200 atlas [Craddock_atlas], with each subject represented by a $200\times 200$ functional connectivity matrix. The BOLD time-series for each ROI are truncated to a length of 100.
@@
The Philadelphia Neuroimaging Cohort (PNC) [pnc] originates from the Brain Behavior Laboratory at the University of Pennsylvania and the Children’s Hospital of Philadelphia. We preprocess this rs-fMRI dataset through a series of steps including slice timing correction, motion correction, registration, normalization, removal of linear trends, bandpass filtering (0.01-0.1 Hz), and spatial smoothing. After preprocessing and quality control, the dataset includes 504 subjects, with 289 (57.34%) being female. The prediction task involves binary classification for gender prediction. The regions of interest are parcellated using the 264-node atlas defined by Power et al. [power2011functional]. Each subject has time-series data collected through 120 time steps.
@@
The Human Connectome Project (HCP) [hcp] is a systematic effort to map macroscopic human brain circuits and their relationship to behavior in a large population of healthy adults. For our analysis, we utilize resting-state fMRI data from the HCP dataset, processed using the standard pipeline in the CONN toolbox [conn], involving realignment, co-registration, normalization, and smoothing of the raw images. Confounding effects from motion artifacts, white matter, and cerebrospinal fluid (CSF) are regressed out of the signal. Subsequently, the BOLD time series are band-pass filtered (0.01-0.1 Hz) to retain relevant functional connectivity information. We focus on the 132 nodes defined on the Harvard-Oxford atlas [desikan2006automated]. After rigorous quality control, this dataset includes 982 samples. Each sample is represented by a $132\times 132$ functional connectivity matrix. The prediction task is fluid intelligence score prediction, treated as a regression problem, with the scores having a mean of 17.03 and a standard deviation of 4.70. To ensure consistency, the BOLD time-series are truncated to a length of 1024.
@@
The Adolescent Brain Cognitive Development Study (ABCD) [abcd] is a large-scale, longitudinal project that recruits children aged 9-10 years across 21 sites in the United States. This study includes repeated imaging scans, along with extensive psychological and cognitive assessments. For our analysis, we use rs-fMRI data from the baseline visit, which is processed using the standard, open-source ABCD-HCP BIDS fMRI pipeline [^1] https://github.com/DCAN-Labs/abcd-hcp-pipeline. The preprocessing pipeline includes multiple stages, starting with distortion correction, anatomical alignment, and surface reconstruction. Further, the functional images undergo motion correction, distortion removal and registration. ANTs-based bias field correction and global signal regression are used to enhance data quality. The time series are band-pass filtered between 0.008 and 0.09 Hz. After quality control, we get 7,327 samples, each represented by a $360\times 360$ connectivity matrix, constructed by calculating the Pearson correlation between the BOLD time-series of each pair of ROIs. The regions of interest (ROIs) are defined based on the HCP 360 atlas [hcp360]. The prediction task is fluid intelligence score prediction with a mean of 18.21 and a standard deviation of 3.69, treated as a regression problem. Considering the variability in sequence lengths within the ABCD dataset, the BOLD time-series are truncated to a unified length of 512 for consistency.
We construct functional connectivity (FC) using Pearson correlation between regional fMRI time series, following the predominant practice in fMRI-based connectomics research [braingb, bnt, ibgnn, neurograph, braingnn, pos_neg3, finn2015functional, cmep]. This choice maintains methodological continuity with prior work and facilitates direct comparison and interpretation by the broader neuroimaging community.
We intentionally select four fMRI datasets with distinct acquisition protocols, preprocessing pipelines, parcellation schemes, sample sizes and prediction tasks to evaluate the robustness and generalizability of our findings. All analyses are performed independently within each dataset, and conclusions are drawn based on consistent relative model performance patterns rather than absolute cross-dataset comparisons. This design ensures that the observed trends are not tied to a specific preprocessing pipeline or participant population, but instead reflect a more general property of GDL models in functional brain connectome analysis.
Benchmarked baselines
@@
For the classical machine learning baselines, we include Connectome-based Predictive Modeling (CPM) [shen2017using], a widely recognized model in network neuroscience for predicting behavioral or clinical outcomes based on brain connectivity data. CPM is typically split into two variations: CPM-NEG and CPM-POS, where predictive features are selected based on negative and positive correlations with the target variable, respectively. Covariance Maximizing Eigenvector-based Predictive Modeling (CMEP) [cmep], a more recent model that first generates candidate features based on eigenvectors with strong linear covariance with the target variable. These features are then used in ElasticNet regression to assess their predictive power, streamlining the process by avoiding arbitrary feature selection thresholds. Additionally, we employed classical models including Logistic Regression for classification, Linear Regression for regression, ElasticNet for both classification and regression, Support Vector Machine (SVM) for classification, Support Vector Regression (SVR) for regression, Random Forest for both classification and regression, as well as Naive Bayes for classification and Kernel Ridge Regression for regression. An overview of these models can be found in [hastie2009elements].
@@
These models are chosen for their established use in brain connectome analysis, ensuring that our findings could be directly compared and validated against existing literature. For all models except CMEP, the input features are the vectorized upper triangular part of the functional brain network matrix, a setting commonly used in classical ML-based functional brain network analysis studies [pos_neg3, pos_neg2, finn2015functional, shen2017using, cui2018effect, PERVAIZ2020116604]. CMEP uses the entire symmetric functional brain network matrix as input.
@@
Most of these classical ML models require a feature selection process. For CPM, feature selection follows the original methodology by applying a p-value threshold [shen2017using] to the correlation between each connection (feature) and the target variable. Features with p-values below the threshold are retained for model training and prediction. CMEP, however, does not require explicit feature selection as it inherently identifies predictive features during its eigenvector generation process. For the other classical ML models, feature selection involves selecting the top $m$ features with the smallest p-values. Specifically, for classification tasks (e.g., using the ABIDE and PNC datasets), p-values are determined through a t-test, while for regression tasks, they are calculated based on the correlation between features and the target variable.
@@
For the non-graph deep learning baselines, we adopt BrainNetCNN [brainnetcnn], a convolutional neural network (CNN) framework specifically designed for predicting clinical neurodevelopmental outcomes from brain networks. Unlike classical image-based CNNs that use spatially local convolutions, BrainNetCNN incorporates novel convolutional filters that operate on the topological locality of structural brain networks. These filters include edge-to-edge, edge-to-node, and node-to-graph convolutions, making the model particularly well-suited for leveraging the unique structure of brain connectivity data. Additionally, we also use two multi-layer perceptron (MLP)-based models: MLP-Flatten and MLP-Node. MLP-Flatten takes the upper triangular part of the brain connectome matrix as input, flattens it into a feature vector, and processes it through two fully connected layers. In contrast, MLP-Node transforms each node’s feature individually using shared weights across nodes, then concatenates these transformed node features into a graph-level embedding for the final prediction.
For the aggregation-based graph deep learning baselines, we focus on models that rely on message aggregation, a core mechanism through which nodes integrate information from their neighbors to update their representations. This process is critical for capturing both structural and relational properties in graph data. Our study includes general-purpose graph deep learning architectures as well as models specifically adapted to the characteristics of brain connectomes.
@@
The general-purpose GDL baselines include following models: Graph Convolutional Network (GCN) [gcn], a fundamental graph neural network (GNN) model that generalizes convolutional operations to graph structures. It aggregates the features of neighboring nodes by applying a convolution-like operation, which computes a weighted sum of the neighboring node features. Graph Attention Network (GAT) [gat] introduces an attention mechanism, allowing the model to assign different importance to neighboring nodes. The attention coefficients are adaptively learned during training, enabling the network to focus on the most relevant neighbors. This adaptive aggregation strategy makes GAT particularly useful for capturing the heterogeneity in node connections, where certain connections might be more relevant to the task. Graph Isomorphism Network (GIN) [gin] is designed to be as powerful as the Weisfeiler-Lehman graph isomorphism test. It aggregates neighboring node features by applying a multi-layer perceptron (MLP) to the sum of the features. GraphSAGE [graphsgae] updates node embeddings by averaging the features of neighboring nodes and passing the result through an MLP.